# Bringing NeRFs to the Latent Space: Inverse Graphics Autoencoder

### v1

@@ 5.5 Ablations
To justify our design choices, we conduct an ablation study in which we alternatively omit our 3D regularization pipeline and our AE preservation components. Ablation results on held-out Objaverse scenes for Vanilla NeRF can be found in Table 2. “IG-AE (no Pr)” removes the AE preservation components from our training by deactivating $\mathfrak{L}_{\mathrm{ae}}$. “IG-AE (no 3D)” omits our 3D regularizatiom by deactivating $\mathfrak{L}_{\mathrm{3D}}$. As illustrated, IG-AE (no Pr) presents deteriorated autoencoding performance due to overfitting on the task of 3D-regularization. IG-AE (no 3D) showcases good autoencoding performances, but lower NVS performances, as the latent space here is not 3D-aware due to the lack of 3D-regularization. Our method (IG-AE) demonstrates both good NVS and autoencoding reconstruction performances as compared to its ablations, thereby justifying the components of our pipeline. Note that ablating both our 3D regularization components and our AE preservation components leads back to the case of a standard autoencoder (AE). Additional ablations done with other NeRF models can be found in Table 4 in Section B.2.
Table 2: Ablations. A comparison between our proposed IG-AE and the baseline autoencoder (AE) on two tasks. First, we compare their NVS performances when used to train a latent Vanilla-NeRF. Second, we compare their reconstruction performances when auto-encoding held-out images from ImageNet. Our method presents both strong NVS and reconstruction performances as compared to its ablations, thus validating our design choices. The bold and underlined entries respectively indicate the best and second-best results.
@@ 6 Conclusion
In this paper, we introduce 3D-aware latent spaces and propose IG-AE, an autoencoder compatible with latent NeRF training. Moreover, we present a latent NeRF training pipeline that brings NeRF architectures to the latent space. We integrate our pipeline into an open-source extension of Nerfstudio, thereby enabling latent NeRF training for its supported architectures. Extensive experiments show the notable improvements our 3D-aware latent space brings as compared to training NeRFs in a standard latent space, as well as the efficiency improvements it brings with respect to training NeRFs in the RGB space. In concluding this paper, several directions of future work arise, as we consider this work to the first milestone towards foundation inverse graphics autoencoders. This includes, for instance, the exploration of approaches improving the representation of high-frequency details when decoding latent NeRFs. We hope that our proposed Nerfstudio extension, as well as our open-source codebase, will promote further research in this direction. In this regard, Kassab et al. (2024) propose a novel method tackling “scaled inverse graphics”, in which our 3D-aware latent space serves as a key component.
Appendix A Tri-Plane Representations
@@
Appendix B Supplementary Results
B.1 RGB Alignment Value
@@
B.2 Quantitative Results
Table 4 includes additional results on held-out scenes from Objaverse. Table 5 compares IG-AE training and RGB training of Nerfstudio methods, in terms of NVS performance as well as training and rendering times.
B.3 Qualitative Results
Figs. 6, 7 and 8 compares the renderings of NeRFs respectively trained on the Cake, House and Figurine scene from Objaverse using our ablated models as well as our full model. Latent NeRFs trained with IG-AE (no 3D) present artifacts when decoding their renderings to the RGB space, as the latent space is not 3D-aware. IG-AE (no Pr) and IG-AE both have 3D-aware latent space and present no artifacts, as previously illustrated, IG-AE (no Pr) demonstrates degraded auto-encoding performances.
Fig. 9 illustrates real images that are auto-encoded using all our models.
Appendix C Hyperparameters
This section is dedicated to illustrate our hyperparameter settings for both our IG-AE training and our Latent NeRF training in Nerfstudio.
C.1 IG-AE Training Settings
@@
C.2 Nerfstudio Training Settings
@@
Table 4: Results on Objaverse scenes. All results are obtained by training the NeRF models using our Latent NeRF Training Pipeline on held-out Objaverse scenes. The IG-AE latent space is more suited for latent NeRF training compared to a standard latent space. This table also extends Table 2 by conducting our ablations on all our adopted representations.

### v2

@@ 5.3 Ablations
To justify our design choices, we conduct an ablation study in which we alternatively omit our 3D regularization pipeline and our AE preservation components. Ablation results on held-out Objaverse scenes for Vanilla NeRF can be found in Table 2. “IG-AE (no Pr)” removes the AE preservation components from our training by deactivating $\mathfrak{L}_{\mathrm{ae}}$. “IG-AE (no 3D)” omits our 3D regularizatiom by deactivating $\mathfrak{L}_{\mathrm{3D}}$. As illustrated, IG-AE (no Pr) presents deteriorated autoencoding performance due to overfitting on the task of 3D-regularization. IG-AE (no 3D) showcases good autoencoding performances, but lower NVS performances, as the latent space here is not 3D-aware due to the lack of 3D-regularization. Our method (IG-AE) demonstrates both good NVS and autoencoding reconstruction performances as compared to its ablations, thereby justifying the components of our pipeline. Note that ablating both our 3D regularization components and our AE preservation components leads back to the case of a standard autoencoder (AE). Additional ablations done with other NeRF models can be found in Table 5 in Section D.2.
Table 2: Ablations. Comparison between IG-AE and the baseline AE on two tasks. First, we compare their NVS performances when used to train a latent Vanilla-NeRF. Second, we compare their reconstruction performances when auto-encoding held-out images from ImageNet. Our method presents both strong NVS and reconstruction performances as compared to its ablations, validating our design choices. The bold and underlined entries indicate the best and second-best results.
@@
5.4 Limitation
Table 8 compares latent NeRF training with classical RGB training. Despite speedups both in terms of training and rendering times, latent NeRFs have a lower NVS performance. This is due to a limitation in the representation of high frequencies in latent NeRFs. We speculate that this issue arises because enforcing 3D consistency in the latent space tends to prioritize regularizing low-frequency structures over high-frequency details. This leads to the loss of high-frequency details when decoding latent NeRF renderings back to the RGB space, as it can be seen in Fig. 5. Furthermore, this currently hinders the applicability of latent NeRFs on complex scenes with abundant high-frequency details, such as real-world scenes. We leave this to future work, as our approach is the first to propose a 3D-aware latent space via an image autoencoder, and is compatible with various NeRF architectures.
@@ 6 Conclusion
In this paper, we propose IG-AE, the first image autoencoder embedding a 3D-aware latent space. Moreover, we present a latent NeRF training pipeline that brings NeRF architectures to the latent space. We integrate our pipeline into an open-source extension of Nerfstudio, thereby enabling latent NeRF training for its supported architectures. Extensive experiments show the notable improvements our 3D-aware latent space brings as compared to training NeRFs in a standard latent space, as well as the efficiency improvements it brings with respect to training NeRFs in the RGB space. In concluding this paper, several directions of future work arise, as we consider this work to be the first milestone towards foundation inverse graphics autoencoders. This includes, for instance, the exploration of approaches improving the representation of high-frequency details when decoding latent NeRFs. We hope that our proposed Nerfstudio extension, as well as our open-source codebase, will promote further research in this direction.
Appendix A Related Work
Table 3: Comparison with related work. Our work is the first to train various NeRF architectures in the latent space of an autoencoder, where NeRFs are able to render images that match AE latent, thanks to the 3D-aware latent space of our IG-AE.
Method Rendering space 3D alignment of rendering and reference spaces NeRF Repr. Task NeRF-VAE RGB yes (RGB) Vanilla-NeRF Generation Rodin RGB yes (RGB) Tri-Planes Generation LN3Diff RGB yes (RGB) Tri-Planes Generation StyleNeRF Feature Vanilla-NeRF Generation Latent-NeRF VAE Latent (Train) RGB (Inference) yes (RGB) I-NGP Generation LERF RGB & CLIP yes (RGB) Vanilla-NeRF 3D Distillation N3F RGB & DINO no (DINO) Tri-Planes 3D Distillation Latent-Editor VAE Latent no (VAE) Vanilla-NeRF 3D Editing ED-NeRF VAE Latent no (VAE) TensoRF 3D Editing RLS RGB & VAE Latent (Train) VAE Latent (Inference) no (VAE) Vanilla-NeRF Reconstruction Ours IG-AE Latent (3D-aware) yes (IG-AE) Vanilla-NeRF, K-Planes, TensoRF, Instant-NGP Reconstruction
This section is dedicated to compare some recent works utilizing NeRFs with scene embeddings and feature images. Table 3 summarizes this section.
As discussed in the related work (Section 2), some methods encode scene information (e.g. images, text) into a scene embedding that then serves to produce a NeRF. Such works do not tackle changing the RGB rendering space of NeRFs.
  • NeRF-VAE (Kosiorek et al., 2021) trains a VAE to encode scene images into a scene embedding. This embedding is then used to condition a NeRF model. They can additionally sample such embeddings in the VAE latent space, which allows for scene generation.
  • Rodin (Wang et al., 2023) enables scene generation with an optional conditioning via images and/or text. They encode the input image and/or textual prompt into a scene embedding, which is then used to condition a diffusion model that generates a Tri-Plane representing the scene.
  • LN3Diff (Lan et al., 2024) enables fast text/single-image-to-3D generation. For single-image-to-3D, the image is encoded into a scene embedding, which is used to condition a DiT transformer that generates a Tri-Plane representing the scene.
Other works do change the rendering space of NeRFs, which makes them more closely related to our work, with still some key differences.
  • StyleNeRF (Gu et al., 2022) proposes a high-resolution generative model with multi-view consistency. They achieve style-based generation via NeRFs that render feature images followed by a decoder. However, these features can only be obtained via the NeRF rendering procedure.
  • Latent-NeRF (Metzer et al., 2023) tackles text-to-3D object generation. It trains latent NeRFs such that their renderings match the posterior distribution of Stable Diffusion (Rombach et al., 2022) under a descriptive text prompt. After this optimization, the latent NeRF is converted into an RGB NeRF to resolve the issues arising from training NeRFs in a standard latent space.
  • LERF (Kerr et al., 2023) distills CLIP features in NeRFs while simultaneously learning the RGB components of the scene, which allows for language queries in 3D.
  • N3F (Tschernezki et al., 2022) distills DINO features into a NeRF representation, while simultaneously learning the RGB components of the scene. This allows to apply image-level tasks (e.g. scene editing, object segmentation) on NeRF renderings. The DINO features rendered by NeRFs represent a 3D-consistent version of the original DINO features.
  • Latent-Editor (Khalid et al., 2023) and ED-NeRF (Park et al., 2024) achieve scene editing by training NeRFs in a VAE latent space. To make this possible, they implement special layers (“adapter” and “refinement layer”) that correct the NeRF renderings to match latent images.
  • RLS (Aumentado-Armstrong et al., 2023) is the closest to our work. It employs hybrid NeRFs that are trained to simultaneously render both RGB and latent components of a scene, where the latent components are learned via a VAE decoder. At inference, it exclusively renders the latent components. However, due to the lack of a 3D-aware latent space, it still requires RGB components at training time to supervise the geometry of the scene.
The presented methods either tackle a different task or use a different rendering space, which makes them not directly comparable with our work. This means that our work is the first to propose training purely latent NeRFs, where no RGB components are utilized, for the reconstructive task.
Appendix B Implementation Details
B.1 Nerfstudio Extension
We extend Nerfstudio to support latent scene learning. Specifically, we incorporate our latent NeRF training pipeline as well as IG-AE in Nerfstudio while adhering to its coding conventions, thereby enabling the training of any Nerfstudio model in our 3D-aware latent space. We modify the ray batching process to make it correspond to full images, where the default behavior uses randomly sampled pixels. This is necessary to be able to decode the rendered images without breaking the gradient flow. Beyond reproducibility purposes for this paper, we hope the proposed generic extension will facilitate research on latent NeRFs and their applications. Our code is open-source and available on the following GitHub repository: https://github.com/AntoineSchnepf/latent-nerfstudio .
B.2 Nerfstudio Models Training Details
We test the following models from Nerfstudio: Vanilla-NeRF (Mildenhall et al., 2020), Instant-NGP (Müller et al., 2022), TensoRF (Chen et al., 2022), and K-Planes (Fridovich-Keil et al., 2023). For each method, the Nerfstudio framework provides a proprietary loss $F_{\theta}$, as well as a custom training procedure that includes specific optimizers, schedulers and other method-specific components. To train a latent NeRF in Nerfstudio, we first train the chosen model for 10 0001000010\,00010 000 iterations to minimize $\mathfrak{L}_{\mathrm{LS}}$ using the method-specific optimization process. Subsequently, we continue the training with 15 0001500015\,00015 000 iterations of RGB alignment by minimizing $\mathfrak{L}_{\mathrm{align}}$. To account for the change of image representations, we modulate the learning rate of each method by a factor of $\xi_{\mathrm{LS}}$ in latent supervision, and a factor $\xi_{\mathrm{align}}$ for RGB alignment. Section F.2 details the hyper-parameters we used in Nerfstudio, including the values of these factors for each method.
B.3 IG-AE Training Details
We adopt the pre-trained “Ostris KL-f8-d16” VAE (Burkett, 2024) from Hugging Face, which has a downscale factor $l=8$, and $c=16$ feature channels in the latent space. We apply a Total Variation (TV) regularization on our Tri-Planes with a factor $\lambda_{\mathrm{TV}}^{\mathrm{3D}}$ which prevents overfitting on train views (we exclude it from Eq. 8 for clarity). Analogously, we also apply TV regularization on encoded real images $k_{j}$. A detailed presentation of TV and its benefits for NVS can be found in Appendix C. Training IG-AE takes $60$ hours on 4×4$\times$4 × NVIDIA L4 GPUs. Our detailed hyper-parameter settings can be found in Section F.1. The training code for IG-AE is open-source and available on the following GitHub repository: https://github.com/k-kassab/igae .
Appendix C Tri-Plane Representations
@@
Appendix D Supplementary Results
D.1 Importance of RGB Alignment
@@
D.2 Additional Quantitative Results
Tables 5, 10, 11, 12, 13 and 14 illustrate additional quantitative evaluations of various NeRF architectures trained in a standard AE latent space and our IG-AE 3D-aware latent space on held-out Objaverse scenes, using our latent NeRF training pipeline. These results further support our conclusions. Additionally, Table 5 also includes our ablations.
Table 6 compares the NVS performance of latent NeRF methods on latent images. Specifically, it compares NeRF-rendered latent images with encoded latent images. Methods trained in the latent space of our IG-AE exhibit significantly better latent NVS performance, confirming its 3D-awareness.
Table 7 compares the autoencoding performance of our IG-AE and the baseline AE when used on images from held-out Objaverse scenes and OOD Shapenet scenes. Our IG-AE showcases better PSNR and SSIM metrics but slightly worse LPIPS metrics, which indicates a minor loss in high frequency details.
Table 8 compares RGB NeRFs and Latent NeRFs using IG-AE, on various Nerfstudio methods, in terms of NVS performance as well as training and rendering times.
D.3 Additional Qualitative Results
Figs. 6, 7 and 8 compares the renderings of NeRFs respectively trained on the Cake, House and Figurine scene from Objaverse using our ablated models as well as our full model. Latent NeRFs trained with IG-AE (no 3D) present artifacts when decoding their renderings to the RGB space, as the latent space is not 3D-aware. IG-AE (no Pr) and IG-AE both have 3D-aware latent space and present no artifacts, as previously illustrated. IG-AE (no Pr) demonstrates degraded auto-encoding performances.
Fig. 9 illustrates real images that are auto-encoded using all IG-AE ablations.
Figs. 10, 11, 12, 13 and 14 showcase additional qualitative results on held-out objaverse scenes.
Novel view synthesis consistency is better visualized in video format, which showcases frame-to-frame consistency. We provide such visualizations on our project page: https://ig-ae.github.io .
Appendix E Choice of autoencoder
Table 9 compares autoencoders with various latent resolutions. For this comparison, we adopt the autoencoders “kl-f4”, “kl-f8”, “kl-f16” from (Rombach et al., 2022), exhibiting downscale factors of 4, 8, and 16 respectively. Note that the compared autoencoders are not trained into inverse graphics autoencoders to conduct this study. As the table presents, NVS performances are directly correlated with latent resolution. We speculate that this comes from the fact that a higher resolution latent space encourages a more local dependency between the RGB image and its latent representation, which is more advantageous in the context of 3D-awareness. As a baseline for IG-AE, we select an autoencoder with a downscale factor of 8 as a compromise between latent resolution and rendering speed.
Appendix F Hyperparameters
This section details our hyperparameter settings for both our IG-AE training and our Latent NeRF training in Nerfstudio.
F.1 IG-AE Training Settings
@@
F.2 Nerfstudio Training Settings
@@
Table 5: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse. This table extends Table 2 by conducting our ablations on all our adopted representations.
@@
Table 6: Latent NVS performance. Quantitative evaluation of NVS performance of NeRF methods on latent images. The metrics compare NeRF-rendered latent images with encoded latent images. Methods trained in the latent space of our IG-AE exhibit significantly better latent NVS performance, confirming its 3D-awareness.
Objaverse (cake) Shapenet (vase)
PSNR SSIM PSNR SSIM
Vanilla-NeRF AE 38.36 0.854 38.48 0.841
IG-AE 50.31 0.989 48.20 0.979
Instant-NGP AE 35.87 0.768 35.94 0.734
IG-AE 44.11 0.967 43.84 0.955
TensoRF AE 36.79 0.817 36.56 0.780
IG-AE 47.37 0.983 46.34 0.967
K-Planes AE 36.85 0.798 36.53 0.755
IG-AE 44.74 0.969 44.26 0.949
Table 7: Synthetic reconstruction. Comparison of autoencoding performances of IG-AE and the baseline autoencoder (AE) on images from Objaverse held-out scenes and Shapenet OOD scenes. In each dataset, the metrics are averaged over the images of all the test scenes.
Objaverse Shapenet
PSNR SSIM LPIPS PSNR SSIM LPIPS
AE 36.47 0.975 0.005 34.78 0.985 0.007
IG-AE 37.92 0.980 0.011 37.16 0.989 0.009
@@
Table 9: Autoencoder choice. Comparison of the performance of autoencoders with various downscale factors on the task of latent NeRF training. We choose an autoencoder with an intermediate downscale factor of 8 as a compromise between latent NVS performance and training time.
kl-f16 kl-f8 kl-f4
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Vanilla-NeRF 29.58 0.941 0.047 32.22 0.962 0.027 35.38 0.978 0.013
Instant-NGP 23.54 0.844 0.117 26.16 0.900 0.084 30.03 0.948 0.036
TensoRF 25.91 0.891 0.080 28.46 0.934 0.042 33.56 0.972 0.014
K-Planes 27.59 0.917 0.060 30.06 0.943 0.032 33.42 0.966 0.017
@@
Table 10: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse.
Helmet Book Burger PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Vanilla-NeRF AE 32.42 0.973 0.024 30.55 0.967 0.044 33.21 0.968 0.037 IG-AE 33.24 0.977 0.021 32.02 0.973 0.039 34.23 0.972 0.037 Instant-NGP AE 24.63 0.918 0.087 23.61 0.904 0.144 23.52 0.862 0.123 IG-AE 28.27 0.951 0.035 27.22 0.950 0.058 27.48 0.922 0.044 TensoRF AE 28.45 0.954 0.040 26.79 0.947 0.070 27.55 0.931 0.038 IG-AE 32.27 0.972 0.021 30.36 0.968 0.041 31.70 0.962 0.022 K-Planes AE 31.14 0.966 0.027 28.90 0.959 0.051 31.91 0.958 0.040 IG-AE 33.27 0.975 0.019 31.42 0.970 0.039 33.53 0.968 0.037
Figure 10: Supplementary evaluations. Qualitative NVS results on held-out scenes from Objaverse.
Table 11: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse.
Cartoon Dragon Flower PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Vanilla-NeRF AE 29.54 0.974 0.017 32.20 0.981 0.022 35.59 0.984 0.014 IG-AE 30.99 0.981 0.012 33.99 0.987 0.015 37.62 0.989 0.010 Instant-NGP AE 22.58 0.906 0.104 25.32 0.932 0.108 27.66 0.925 0.091 IG-AE 25.68 0.949 0.038 28.42 0.961 0.044 31.07 0.963 0.035 TensoRF AE 25.44 0.946 0.045 28.53 0.964 0.044 30.62 0.962 0.038 IG-AE 29.29 0.973 0.017 32.65 0.983 0.017 35.44 0.985 0.013 K-Planes AE 28.22 0.967 0.021 31.09 0.976 0.027 33.77 0.978 0.023 IG-AE 30.57 0.979 0.012 33.24 0.984 0.016 37.26 0.988 0.011
Figure 11: Supplementary evaluations. Qualitative NVS results on held-out scenes from Objaverse.
Table 12: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse.
Gravestone Hut Painting PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Vanilla-NeRF AE 31.64 0.973 0.034 30.33 0.963 0.030 36.07 0.984 0.018 IG-AE 32.63 0.978 0.029 30.70 0.964 0.027 37.79 0.989 0.013 Instant-NGP AE 22.95 0.885 0.169 22.64 0.881 0.154 27.02 0.930 0.089 IG-AE 26.68 0.941 0.073 26.04 0.933 0.066 34.34 0.980 0.020 TensoRF AE 26.65 0.944 0.076 26.89 0.943 0.058 31.51 0.967 0.037 IG-AE 31.15 0.974 0.032 30.00 0.963 0.030 37.75 0.989 0.012 K-Planes AE 28.72 0.960 0.053 28.68 0.955 0.038 34.79 0.982 0.020 IG-AE 31.25 0.973 0.031 30.19 0.961 0.029 38.28 0.990 0.012
Figure 12: Supplementary evaluations. Qualitative NVS results on held-out scenes from Objaverse.
Table 13: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse.
Pig Ringer Robot PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Vanilla-NeRF AE 33.21 0.981 0.029 32.14 0.975 0.024 29.94 0.962 0.028 IG-AE 34.34 0.987 0.019 33.70 0.982 0.015 31.76 0.973 0.019 Instant-NGP AE 24.61 0.913 0.140 24.68 0.901 0.119 23.49 0.884 0.107 IG-AE 28.35 0.956 0.067 28.24 0.951 0.042 26.01 0.923 0.049 TensoRF AE 28.61 0.959 0.060 27.71 0.947 0.057 26.43 0.933 0.042 IG-AE 32.58 0.981 0.027 32.03 0.976 0.019 29.96 0.963 0.024 K-Planes AE 31.03 0.972 0.044 30.61 0.964 0.032 29.08 0.954 0.031 IG-AE 33.12 0.981 0.026 32.99 0.975 0.020 31.10 0.967 0.022
Figure 13: Supplementary evaluations. Qualitative NVS results on held-out scenes from Objaverse.
Table 14: Supplementary evaluations. Quantitative NVS results on held-out scenes from Objaverse.
Coffin Table Villa PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Vanilla-NeRF AE 30.61 0.968 0.042 33.06 0.978 0.024 30.72 0.970 0.033 IG-AE 33.17 0.976 0.032 34.96 0.983 0.019 32.18 0.977 0.024 Instant-NGP AE 21.67 0.876 0.211 23.29 0.873 0.161 24.21 0.903 0.138 IG-AE 27.25 0.941 0.071 29.19 0.956 0.042 26.18 0.935 0.084 TensoRF AE 27.35 0.946 0.076 27.16 0.943 0.071 26.86 0.943 0.065 IG-AE 31.87 0.973 0.034 33.20 0.979 0.020 30.59 0.971 0.031 K-Planes AE 29.30 0.960 0.053 30.72 0.965 0.036 28.86 0.957 0.048 IG-AE 31.56 0.971 0.036 33.25 0.976 0.023 31.66 0.972 0.027
Figure 14: Supplementary evaluations. Qualitative NVS results on held-out scenes from Objaverse.